# Deep Submodular Peripteral Networks

### 13 Mar 2024 v1

@@ 1 Introduction and Background
The first is that of identifying a useful practical scalable submodular function that can be used for real-world data science tasks. Submodular functions have received considerable attention in the field of machine learning, leading to the creation of algorithms that offer near-optimal solutions for various applications. These applications include detecting disease outbreaks Leskovec et al. (2007), modeling fine structure in computer vision Jegelka and Bilmes (2011), summarizing images Tschiatschek et al. (2014); Mitrovic et al. (2018); Feldman et al. (2018); Harshaw et al. (2022), active learning Guillory and Bilmes (2011); Golovin and Krause (2010); Esfandiari et al. (2021), compressed sensing, structured convex norms, and sparsity Bach (2010); Elenberg et al. (2018), fairness Celis et al. (2016); Kazemi et al. (2021), efficient model training Lin and Bilmes (2009, 2010); Mirzasoleiman et al. (2016, 2020) recommendation systems Mitrovic et al. (2019), causal structure Zhou and Spanos (2016); Sussex et al. (2021), and brain parcellating Salehi et al. (2017) (see also the reviews Tohidi et al. (2020); Bilmes (2022)). Despite these myriad applications, most research on submodularity has been on the algorithmic and theoretical side which assumes the submodular function needing to be optimized is already at hand. While there has been work showing the theoretical challenges associated with learning submodularity Balcan et al. (2011, 2016); Balcan and Harvey (2018), there is relatively little work on practical methods to produce a useful submodular function in the first place (exceptions include Lin and Bilmes (2012); Sipos et al. (2011); Tschiatschek et al. (2018)). Often, because it works well, one uses the non-parametric submodular facility location (FL) function, i.e., for $A\subseteq V$ with $|V|=n$, $f(A)=\sum_{i=1}^{n}\max_{a\in A}s_{a,i}$ where $s_{a,i}\geq 0$ is a similarity between items $a$ and $i$. On the other hand, the FL function’s computational and memory costs grow quadratically with dataset size $n$ and so FL is infeasible for large data or online streaming applications. There is a need for practical strategies to obtain useful, scalable, general-purpose, and widely applicable submodular functions.
@@
A special case of such preference elicitation from human oracles have been studied going back many decades. For example, the psychometric “Law of comparative judgment" Thurstone (1927) focuses on establishing numeric interval scales based on knowing preferences between pairs of choices (e.g., $A$ vs $B$). The Bradley-Terry Bradley and Terry (1952) model (generalized to ranking elicitations in the Luce-Shephard model Luce (1959); Shepard (1957)) also involves preferences among elements of pairs. These preferences, however, are not graded and rather are either only binary ($A\prec B$ or $B\prec A$), ternary (allowing also for indifference), or quaternary (allowing also for incomparability). Thus, each oracle query provides minimal information about a pair. Therefore, multiple preference ratings are often aggregated where a collection of (presumed random i.i.d.) oracles vote on the preference between items of the same pair. This produces a histogram of preference ratings $C_{ij}$ (a count of how many times $i$ is preferred to $j$ Silverstein and Farrell (2001)) and this can be used in modern reinforcement learning with human feedback (RLHF) systems Wirth et al. (2017); Christiano et al. (2023); Ouyang et al. (2022); Zhan et al. (2023); Wang et al. (2023) for reward learning (i.e., inverse reinforcement learning Arora and Doshi (2021); Adams et al. (2022)). Overall, the above modeling strategies are known in the field of psychometric theory Guilford (1954) as “paired comparisons” where it is said that “paired comparison methods generally give much more reliable results” Nunnally and Bernstein (1994) for models of human preference expression and elicitation.
@@
There are other models of preference elicitation besides paired comparisons, however. Our aforementioned desire is for “graded pairwise comparisons” (GPC) which also have been studied for quite some time Scheffe (1952); Bechtel (1967, 1971). Quite recently, it was found that “GPCs are expected to reduce faking compared with Likert-type scales and to produce more reliable, less ipsative [the extent to which results are influenced by the specific method of measurement] trait scores than traditional-binary forced-choice formats” Lingel et al. (2022) such as paired comparisons. An intuitive reason for this is that, once we get to the point that an oracle is asked only if $A\prec B$ or $B\prec A$, this does not elicit as much information out of the oracle as would asking for $\text{\emph{Score}}({A},{B})$ even though the incremental oracle effort for the latter is often negligible. As far as we know, despite GPC’s potential information extraction efficiency advantages, GPC has not been addressed in the modern machine learning community. Also, while histograms $C_{i,j}$ are appropriate for maximum likelihood estimation (MLE) of non-linear logistic and softmax style regression Christiano et al. (2023); Ouyang et al. (2022); Zhu et al. (2023); Zhan et al. (2023), MLE with a Bradley-Terry style model is suboptimal to learn scoring from graded preferences $\text{\emph{Score}}({A},{B})\in\mathbb{R}$ which can be an arbitrary signed real number.
@@
We are inspired by contrastive learning Balestriero et al. (2023), which has pushed the boundaries of deep learning by leveraging the massive information in unlabeled data. Unlike standard supervised learning which teaches models to associate samples to human-supplied labels, contrastive learning encourages models to capture relationships between pairs, groups, or sections of unlabeled samples. It does this by “contrasting” positive pairs (i.e., similar data points) against negative pairs (dissimilar data points). Mathematically, this approach uses loss functions that encourage the model to learn embeddings where positive pairs become “close” (according to some distance measure), while negative pairs become “far.” A common setting is the contrastive loss, such as the triplet loss or the Siamese network loss. For instance, in the triplet loss framework, the model is trained on triplets of data points consisting of an anchor $x$, a positive example $x^{+}$ (similar to the anchor), and a negative example $x^{-}$ (dissimilar from the anchor). Optimizing a model with such a loss function encourages a representation that respects these positive/negative pairwise preferences.
@@
On the other hand, such contrastive approaches are analogous to binary-only pairwise comparison elicitations. That is, existing contrastive losses neither utilize information regarding how positive the pair $x,x^{+}$ is nor how negative the pair $x,x^{-}$ is. A more nuanced graded comparison, if available, would be an improvement. In fact, the benefit of graded comparisons is analogous to the benefits of using soft over hard labels in supervised training (say from model distillation) which has benefits in uncertainty representation, regularization, noise mitigation, adjustable smoothness via softmax temperature, and more efficient knowledge transfer. Another limitation of previous contrastive learning is that it operates only with positive or negative groups of size two (pairs) rather than larger groups. While it is often the case that the anchor $x$ is contrasted with many negative examples, it is always done so pairwise, considering each $x^{-}$ with $x$ together in a sum of pairwise terms. Also, while Tian et al. (2020) is the only work we are aware of that considers more than two positive examples at a time, each of these positive samples is again considered only paired up with a negative sample, leading to ($k\choose 2$ pairings with $k$ positive examples. Sums of pairwise terms do not allow for the consideration of higher order relationships, e.g., a given pair might be measured differently in the context of a third sample. In addition to addressing the two aforementioned problems, DSPNs and our GPC-style loss addresses this issue as well, although our learning goal in this paper is learning submodularity not in representation learning (the latter of which we leave to future work).
@@
An outline of our paper follows. In §2 we define the class of DSPN functions, motivating the name “peripteral”, and offering theoretical comparisons with the class of deep sets Zaheer et al. (2017). Next, §3 describes a new GPC-style “peripteral” loss that is appropriate to train a DSPN but can be used in any GPC-based preference elicitation learning task. For DSPN learning, this happens by producing a list of pairs of subsets $E,M$ that can be used to transfer information from the oracle target to the DSPN being learnt. §4 describes several $E,M$ pair sampling strategy including an active-learning inspired submodular feedback approach. §5 empirically evaluates DSPN learning based on how effectively information is transferred from the target oracle and performance on downstream tasks such as experimental design. We find that the peripteral loss outperforms other losses for training DSPNs, and show that the DSPN architecture is more effective at learning from the target function than baseline methods such as Deep Sets and SetTransformers.
@@ 2 Deep Submodular Peripteral Networks
The second stage of a DSPN is a submodular preserving and permutation-invariant aggregation. Before describing this generally, we start with a simple example. Each element ${\phi_{{w_{\mathfrak{p}}}}}(x_{a})_{j}$ of ${\phi_{{w_{\mathfrak{p}}}}}(x_{a})$ is a score for $x_{a}$ along dimension $j$ and so a non-negative modular set function can be constructed via $m_{j}(A)=\sum_{a\in A}{\phi_{{w_{\mathfrak{p}}}}}(x_{a})_{j}$. Such a summation is an aggregator that, along dimension $j$ simply sums up the $j^{\text{th}}$ contribution for every object in $A$. A simple way to convert this to a monotone non-decreasing submodular function composes it with a non-decreasing concave function $\psi$ yielding $h_{j}(A)=\psi(m_{j}(A))$ — performing this for all $j$ yields a $d$-dimensional vector of submodular functions. More generally, any aggregation function $\oplus$ that preserves the submodularity of $h_{j}(A)=\psi(\oplus_{a\in A}m_{j}(a))$ would suffice so long as it is also permutation invariant Zaheer et al. (2017) (see Def. 6). For example, with $\oplus=\max$ (i.e., max pooling), the function $\psi(\max_{a\in A}m_{j}(a))$ is again submodular. We show that there is an expressive infinitely large family of submodular preserving permutation-invariant aggregation operators, taking the form of the weighted matroid rank functions (Def. 5) which are defined based on a matroid $\mathcal{M}=(V,\mathcal{I})$ and a non-negative vector $m\in\mathbb{R}_{+}^{|V|}$.
@@
A proof is in Appendix B. We identify the aggregator $\oplus$ as $\oplus_{a\in A}m_{j}(a)=\text{rank}_{\mathcal{M},m_{j}}(A)$ and produce a new submodular function $h_{j}^{w_{\mathfrak{p}}}(A)=\psi(\text{rank}_{\mathcal{M},{\phi_{{w_{\mathfrak{p}}}}}(x_{:})_{j}}(A))$, where the $n$-dimensional vector ${\phi_{{w_{\mathfrak{p}}}}}(x_{:})_{j}$ constitute the weights for the matroid. Depending on the matroid, this yields summation (using a uniform matroid), max-pooling (using a partition matroid), and an unbounded number of others thanks to the flexibility of matroids Oxley (2011). In theory, this means that the aggregator could also be trained via discrete optimization over the space of discrete matroid structures (in §5, however, our aggregators are fixed up front). The second stage of the DSPN thus aggregates the modular outputs of the first stage into a resulting $d$-dimensional vector for any set $A\subseteq V$, namely $h_{w_{\mathfrak{p}}}(A)=(h_{1}^{w_{\mathfrak{p}}}(A),h_{2}^{w_{\mathfrak{p}}}(A),\dots,h_{d}^{w_{\mathfrak{p}}}(A))\in\mathbb{R}_{+}^{d}$. This can also be viewed as a vector of submodular functions. We presume in this work that each aggregator uses the same matroid and each pillar dimension is used with exactly one matroid, but this can be relaxed leading to a $d^{\prime}>d$ dimensional space — in other words, many different discrete matroid structures can be used at the same time if so desired.
The third and final “roof” stage of a DSPN is a deep submodular function (DSF) Dolhansky and Bilmes (2016). Deep submodular functions are a nested series of vectors of monotone non-decreasing concave functions composed with each other, layer-after-layer, leading to a final single scalar output. Each layer contained in a DSF has non-negative weight values which assures the DSF is submodular. It was shown in Bilmes and Bai (2017) that the family of submodular functions expressible by DSFs strictly increase with depth. DSFs are also trainable using gradient-based methods analogous to how any deep neural network can be trained. A DSF is defined as a set function, but we utilize the “root”111”root” here is distinct from ”roof”. Bilmes and Bai (2017) multivariate multi-layered concave function ${\psi_{{w_{\mathfrak{r}}}}}:\mathbb{R}^{d}\to\mathbb{R}_{+}$ of a DSF in this work. The final stage of the DSPN composes the output of the second stage with this final DSF via $f_{w}(A)={\psi_{{w_{\mathfrak{r}}}}}(h_{w_{\mathfrak{p}}}(A))$ where $w=({w_{\mathfrak{p}}},{w_{\mathfrak{r}}})$ constitute all the learnable continuous parameters of the DSPN (assuming the discrete parameters, such as the matroid structure, is fixed). While the roof parameters ${w_{\mathfrak{r}}}$ must be positive, the pillar parameters ${w_{\mathfrak{p}}}$ are free, but the pillar must produce non-negative outputs.
@@
Clearly there is a strong relationship between DSPNs and Deep Sets Zaheer et al. (2017). In fact any DSPNs is a special case of a Deep Set, but there is an important distinction, namely that a DSPN is assuredly submodular. It may seem like this would restrict the family of aggregation functions available but, as shown above, there are in theory an unbounded number of aggregators to study. Training a DSPN also preserves submodularity simply by ensuring ${w_{\mathfrak{r}}}$ remains positive which can be achieved using projected gradient descent (the positive orthant is one of the easiest of the constraints to maintain). A deep set, when trained, can easily lose the submodular property, something we demonstrate in §5, rendering the deep set inapplicable when submodularity is required. Also, as was shown in Bilmes and Bai (2017), DSFs are already quite expressive, the only known limitation being their inability to express certain matroid rank functions. With a DSPN, however, the aggregator functions immediately eliminate this inability. Hence, it is an open theoretical problem if a DSPN, by varying in parameter $w$ and combinatorial space $\mathcal{M}$, can represent all possible monotone non-decreasing submodular functions.

### 16 Mar 2024 v2

@@ 1 Introduction and Background
The first is that of identifying a useful practical scalable submodular function that can be used for real-world data science tasks. Submodular functions have received considerable attention in the field of machine learning, leading to the creation of algorithms that offer near-optimal solutions for various applications. These applications include detecting disease outbreaks (Leskovec et al., 2007), modeling fine structure in computer vision (Jegelka and Bilmes, 2011), summarizing images (Tschiatschek et al., 2014; Mitrovic et al., 2018; Feldman et al., 2018; Harshaw et al., 2022), active learning (Guillory and Bilmes, 2011; Golovin and Krause, 2010; Esfandiari et al., 2021), compressed sensing, structured convex norms, and sparsity (Bach, 2010; Elenberg et al., 2018), fairness (Celis et al., 2016; Kazemi et al., 2021), efficient model training (Lin and Bilmes, 2009, 2010; Mirzasoleiman et al., 2016, 2020) recommendation systems (Mitrovic et al., 2019), causal structure (Zhou and Spanos, 2016; Sussex et al., 2021), and brain parcellating (Salehi et al., 2017) (see also the reviews (Tohidi et al., 2020; Bilmes, 2022)). Despite these myriad applications, most research on submodularity has been on the algorithmic and theoretical side which assumes the submodular function needing to be optimized is already at hand. While there has been work showing the theoretical challenges associated with learning submodularity (Balcan et al., 2011, 2016; Balcan and Harvey, 2018), there is relatively little work on practical methods to produce a useful submodular function in the first place (exceptions include (Lin and Bilmes, 2012; Sipos et al., 2011; Tschiatschek et al., 2018)). Often, because it works well, one uses the non-parametric submodular facility location (FL) function, i.e., for $A\subseteq V$ with $|V|=n$, $f(A)=\sum_{i=1}^{n}\max_{a\in A}s_{a,i}$ where $s_{a,i}\geq 0$ is a similarity between items $a$ and $i$. On the other hand, the FL function’s computational and memory costs grow quadratically with dataset size $n$ and so FL is infeasible for large data or online streaming applications. There is a need for practical strategies to obtain useful, scalable, general-purpose, and widely applicable submodular functions.
@@
A special case of such preference elicitation from human oracles have been studied going back many decades. For example, the psychometric “Law of comparative judgment" (Thurstone, 1927) focuses on establishing numeric interval scales based on knowing preferences between pairs of choices (e.g., $A$ vs $B$). The Bradley-Terry (Bradley and Terry, 1952) model (generalized to ranking elicitations in the Luce-Shephard model (Luce, 1959; Shepard, 1957)) also involves preferences among elements of pairs. These preferences, however, are not graded and rather are either only binary ($A\prec B$ or $B\prec A$), ternary (allowing also for indifference), or quaternary (allowing also for incomparability). Thus, each oracle query provides minimal information about a pair. Therefore, multiple preference ratings are often aggregated where a collection of (presumed random i.i.d.) oracles vote on the preference between items of the same pair. This produces a histogram of preference ratings $C_{ij}$ (a count of how many times $i$ is preferred to $j$ (Silverstein and Farrell, 2001)) and this can be used in modern reinforcement learning with human feedback (RLHF) systems (Wirth et al., 2017; Christiano et al., 2023; Ouyang et al., 2022; Zhan et al., 2023; Wang et al., 2023) for reward learning (i.e., inverse reinforcement learning (Arora and Doshi, 2021; Adams et al., 2022)). Overall, the above modeling strategies are known in the field of psychometric theory (Guilford, 1954) as “paired comparisons” where it is said that “paired comparison methods generally give much more reliable results” (Nunnally and Bernstein, 1994) for models of human preference expression and elicitation.
@@
There are other models of preference elicitation besides paired comparisons, however. Our aforementioned desire is for “graded pairwise comparisons” (GPC) which also have been studied for quite some time (Scheffe, 1952; Bechtel, 1967, 1971). Quite recently, it was found that “GPCs are expected to reduce faking compared with Likert-type scales and to produce more reliable, less ipsative [the extent to which results are influenced by the specific method of measurement] trait scores than traditional-binary forced-choice formats” (Lingel et al., 2022) such as paired comparisons. An intuitive reason for this is that, once we get to the point that an oracle is asked only if $A\prec B$ or $B\prec A$, this does not elicit as much information out of the oracle as would asking for $\text{\emph{Score}}({A},{B})$ even though the incremental oracle effort for the latter is often negligible. As far as we know, despite GPC’s potential information extraction efficiency advantages, GPC has not been addressed in the modern machine learning community. Also, while histograms $C_{i,j}$ are appropriate for maximum likelihood estimation (MLE) of non-linear logistic and softmax style regression (Christiano et al., 2023; Ouyang et al., 2022; Zhu et al., 2023; Zhan et al., 2023), MLE with a Bradley-Terry style model is suboptimal to learn scoring from graded preferences $\text{\emph{Score}}({A},{B})\in\mathbb{R}$ which can be an arbitrary signed real number.
@@
We are inspired by contrastive learning (Balestriero et al., 2023), which has pushed the boundaries of deep learning by leveraging the massive information in unlabeled data. Unlike standard supervised learning which teaches models to associate samples to human-supplied labels, contrastive learning encourages models to capture relationships between pairs, groups, or sections of unlabeled samples. It does this by “contrasting” positive pairs (i.e., similar data points) against negative pairs (dissimilar data points). Mathematically, this approach uses loss functions that encourage the model to learn embeddings where positive pairs become “close” (according to some distance measure), while negative pairs become “far.” A common setting is the contrastive loss, such as the triplet loss or the Siamese network loss. For instance, in the triplet loss framework, the model is trained on triplets of data points consisting of an anchor $x$, a positive example $x^{+}$ (similar to the anchor), and a negative example $x^{-}$ (dissimilar from the anchor). Optimizing a model with such a loss function encourages a representation that respects these positive/negative pairwise preferences.
@@
On the other hand, such contrastive approaches are analogous to binary-only pairwise comparison elicitations. That is, existing contrastive losses neither utilize information regarding how positive the pair $x,x^{+}$ is nor how negative the pair $x,x^{-}$ is. A more nuanced graded comparison, if available, would be an improvement. In fact, the benefit of graded comparisons is analogous to the benefits of using soft over hard labels in supervised training (say from model distillation) which has benefits in uncertainty representation, regularization, noise mitigation, adjustable smoothness via softmax temperature, and more efficient knowledge transfer. Another limitation of previous contrastive learning is that it operates only with positive or negative groups of size two (pairs) rather than larger groups. While it is often the case that the anchor $x$ is contrasted with many negative examples, it is always done so pairwise, considering each $x^{-}$ with $x$ together in a sum of pairwise terms. Also, while (Tian et al., 2020) is the only work we are aware of that considers more than two positive examples at a time, each of these positive samples is again considered only paired up with a negative sample, leading to ($k\choose 2$ pairings with $k$ positive examples. Sums of pairwise terms do not allow for the consideration of higher order relationships, e.g., a given pair might be measured differently in the context of a third sample. In addition to addressing the two aforementioned problems, DSPNs and our GPC-style loss addresses this issue as well, although our learning goal in this paper is learning submodularity not in representation learning (the latter of which we leave to future work).
@@
An outline of our paper follows. In §2 we define the class of DSPN functions, motivating the name “peripteral”, and offering theoretical comparisons with the class of deep sets (Zaheer et al., 2017). Next, §3 describes a new GPC-style “peripteral” loss that is appropriate to train a DSPN but can be used in any GPC-based preference elicitation learning task. For DSPN learning, this happens by producing a list of pairs of subsets $E,M$ that can be used to transfer information from the oracle target to the DSPN being learnt. §4 describes several $E,M$ pair sampling strategy including an active-learning inspired submodular feedback approach. §5 empirically evaluates DSPN learning based on how effectively information is transferred from the target oracle and performance on downstream tasks such as experimental design. We find that the peripteral loss outperforms other losses for training DSPNs, and show that the DSPN architecture is more effective at learning from the target function than baseline methods such as Deep Sets and SetTransformers.
@@ 2 Deep Submodular Peripteral Networks
The second stage of a DSPN is a submodular preserving and permutation-invariant aggregation. Before describing this generally, we start with a simple example. Each element ${\phi_{{w_{\mathfrak{p}}}}}(x_{a})_{j}$ of ${\phi_{{w_{\mathfrak{p}}}}}(x_{a})$ is a score for $x_{a}$ along dimension $j$ and so a non-negative modular set function can be constructed via $m_{j}(A)=\sum_{a\in A}{\phi_{{w_{\mathfrak{p}}}}}(x_{a})_{j}$. Such a summation is an aggregator that, along dimension $j$ simply sums up the $j^{\text{th}}$ contribution for every object in $A$. A simple way to convert this to a monotone non-decreasing submodular function composes it with a non-decreasing concave function $\psi$ yielding $h_{j}(A)=\psi(m_{j}(A))$ — performing this for all $j$ yields a $d$-dimensional vector of submodular functions. More generally, any aggregation function $\oplus$ that preserves the submodularity of $h_{j}(A)=\psi(\oplus_{a\in A}m_{j}(a))$ would suffice so long as it is also permutation invariant (Zaheer et al., 2017) (see Def. 6). For example, with $\oplus=\max$ (i.e., max pooling), the function $\psi(\max_{a\in A}m_{j}(a))$ is again submodular. We show that there is an expressive infinitely large family of submodular preserving permutation-invariant aggregation operators, taking the form of the weighted matroid rank functions (Def. 5) which are defined based on a matroid $\mathcal{M}=(V,\mathcal{I})$ and a non-negative vector $m\in\mathbb{R}_{+}^{|V|}$.
@@
A proof is in Appendix B. We identify the aggregator $\oplus$ as $\oplus_{a\in A}m_{j}(a)=\text{rank}_{\mathcal{M},m_{j}}(A)$ and produce a new submodular function $h_{j}^{w_{\mathfrak{p}}}(A)=\psi(\text{rank}_{\mathcal{M},{\phi_{{w_{\mathfrak{p}}}}}(x_{:})_{j}}(A))$, where the $n$-dimensional vector ${\phi_{{w_{\mathfrak{p}}}}}(x_{:})_{j}$ constitute the weights for the matroid. Depending on the matroid, this yields summation (using a uniform matroid), max-pooling (using a partition matroid), and an unbounded number of others thanks to the flexibility of matroids (Oxley, 2011). In theory, this means that the aggregator could also be trained via discrete optimization over the space of discrete matroid structures (in §5, however, our aggregators are fixed up front). The second stage of the DSPN thus aggregates the modular outputs of the first stage into a resulting $d$-dimensional vector for any set $A\subseteq V$, namely $h_{w_{\mathfrak{p}}}(A)=(h_{1}^{w_{\mathfrak{p}}}(A),h_{2}^{w_{\mathfrak{p}}}(A),\dots,h_{d}^{w_{\mathfrak{p}}}(A))\in\mathbb{R}_{+}^{d}$. This can also be viewed as a vector of submodular functions. We presume in this work that each aggregator uses the same matroid and each pillar dimension is used with exactly one matroid, but this can be relaxed leading to a $d^{\prime}>d$ dimensional space — in other words, many different discrete matroid structures can be used at the same time if so desired.
The third and final “roof” stage of a DSPN is a deep submodular function (DSF) (Dolhansky and Bilmes, 2016). Deep submodular functions are a nested series of vectors of monotone non-decreasing concave functions composed with each other, layer-after-layer, leading to a final single scalar output. Each layer contained in a DSF has non-negative weight values which assures the DSF is submodular. It was shown in (Bilmes and Bai, 2017) that the family of submodular functions expressible by DSFs strictly increase with depth. DSFs are also trainable using gradient-based methods analogous to how any deep neural network can be trained. A DSF is defined as a set function, but we utilize the “root”111”root” here is distinct from ”roof”. (Bilmes and Bai, 2017) multivariate multi-layered concave function ${\psi_{{w_{\mathfrak{r}}}}}:\mathbb{R}^{d}\to\mathbb{R}_{+}$ of a DSF in this work. The final stage of the DSPN composes the output of the second stage with this final DSF via $f_{w}(A)={\psi_{{w_{\mathfrak{r}}}}}(h_{w_{\mathfrak{p}}}(A))$ where $w=({w_{\mathfrak{p}}},{w_{\mathfrak{r}}})$ constitute all the learnable continuous parameters of the DSPN (assuming the discrete parameters, such as the matroid structure, is fixed). While the roof parameters ${w_{\mathfrak{r}}}$ must be positive, the pillar parameters ${w_{\mathfrak{p}}}$ are free, but the pillar must produce non-negative outputs.
@@
Clearly there is a strong relationship between DSPNs and Deep Sets (Zaheer et al., 2017). In fact any DSPNs is a special case of a Deep Set, but there is an important distinction, namely that a DSPN is assuredly submodular. It may seem like this would restrict the family of aggregation functions available but, as shown above, there are in theory an unbounded number of aggregators to study. Training a DSPN also preserves submodularity simply by ensuring ${w_{\mathfrak{r}}}$ remains positive which can be achieved using projected gradient descent (the positive orthant is one of the easiest of the constraints to maintain). A deep set, when trained, can easily lose the submodular property, something we demonstrate in §5, rendering the deep set inapplicable when submodularity is required. Also, as was shown in (Bilmes and Bai, 2017), DSFs are already quite expressive, the only known limitation being their inability to express certain matroid rank functions. With a DSPN, however, the aggregator functions immediately eliminate this inability. Hence, it is an open theoretical problem if a DSPN, by varying in parameter $w$ and combinatorial space $\mathcal{M}$, can represent all possible monotone non-decreasing submodular functions.